### v2

# Sink equilibria and the attractors of learning in games

Abstract
Characterizing the limit behavior — that is, the attractors — of learning dynamics is one of the most fundamental open questions in game theory. In recent work in this front, it was conjectured that the attractors of the replicator dynamic are in one-to-one correspondence with the sink equilibria of the game — the sink strongly connected components of a game’s preference graph —, and it was established that they do stand in at least one-to-many correspondence with them. We make threefold progress on the problem of characterizing attractors. First, we show through a topological construction that the one-to-one conjecture is false. The counterexamples derive from objects called local sources — fixed points which lie within the sink equilibrium yet are locally repelling. Second, we make progress on the attractor characterization problem for two-player games by establishing that the one-to-one conjecture is true when a local property called pseudoconvexity holds. Pseudoconvexity prevents the existence of local sources, and generalizes the existing cases — such as zero-sum games and potential games — where the conjecture was known to be true.
1 Introduction
What are the possible outcomes of a collection of jointly-learning rational agents? This is a fundamental — arguably the fundamental — problem in the study of learning in games, with consequences for machine learning, economics and evolutionary biology. The question has received decades of study by mathematicians, economists and computer scientists (Zeeman, 1980; Milgrom and Roberts, 1991; Hofbauer, 1996; Sandholm, 2010; Papadimitriou and Piliouras, 2019) — and yet, it remains broadly unanswered.
One reason for the lack of progress has been the historical focus on Nash equilibria as the outcome of a game (Myerson, 1997). Over time this approach was found to be problematic (Kleinberg et al., 2011) — not only do standard learning algorithms not converge to Nash equilibria in general games (Benaïm et al., 2012; Milionis et al., 2023), but Nash equilibria are also intractable to compute (Daskalakis et al., 2009). Non-convergence to Nash equilibria has been traditionally viewed as a limitation of game dynamics. In a departure, Papadimitriou and Piliouras (2019) argued the opposite: non-convergence of learning to Nash equilibria could instead be viewed as yet another limitation of the Nash equilibrium concept. Instead, the outcomes of learning — whatever they may be — should be the fundamental object of interest in game theory.
Of course, determining the limit behavior of dynamics in games is a very difficult problem. Even the replicator dynamic (Taylor and Jonker, 1978; Smith and Price, 1973) — the continuous-time analog of the multiplicative weights algorithm (Arora et al., 2012) and the flagship dynamic of evolutionary game theory (Hofbauer and Sigmund, 2003; Sandholm, 2010) — can have extremely complex long-run behaviors (Sato et al., 2002). The problem is further complicated by the emergence of chaos and the sensitive dependence on initial conditions. However, Papadimitriou and Piliouras (2019) suggested a possible path forward; they hypothesized that the behavior of the replicator (and possibly of dynamics in general) can be well-approximated by a simple combinatorial tool, called the preference graph of a game (Biggar and Shames, 2023a, 2024, 2025).
[…]
figs/shapley_square
Figure 1: The preference graph of Shapley’s game (Shapley, 1964). It has a unique sink equilibrium, which is the highlighted 6-cycle.
The preference graph is a graphical representation of the underlying structure of the preferences in a normal-form game (Biggar and Shames, 2023a) (see Figure 1). While it has only recently become the subject of study in its own right, it plays an implicit role in many of milestone results across the history of game theory (Biggar and Shames, 2025). By definition, pure Nash equilibria (PNEs) are sinks in the preference graph. This observation tempts one to generalize PNEs in a direction different from the one that Nash pursued, namely to consider the sink equilibria, the sink strongly connected components of the preference graph, as a novel solution concept, and indeed one that is not incompatible with learning dynamics. Papadimitriou and Piliouras (2019) proposed a dramatic hypothesis: the sink equilibria are good proxies for the attractors of the replicator. We call this the attractor–sink equilibrium correspondence hypothesis.
If true, this hypothesis would be extremely useful for analyzing the outcome of games. Sink equilibria are a simple and easily-computable object which is robust to small changes in the payoff structure of the game. This would finally open a path to predicting strategic interactions between realistic learning agents. On the other hand, finding nontrivial mathematical connections between attractors and sink equilibria is difficult. The first significant progress on this hypothesis appeared in Biggar and Shames (2023b), where it was proved that every attractor of the replicator must contain a sink equilibrium, proving sinks are indeed a sort of ‘skeleton’ for the attractors of the replicator. They strengthened the attractor-sink equilibrium hypothesis to a conjecture:
Conjecture 1.1.
The attractors of the replicator in any game are in one-to-one correspondence with the sink equilibria.
They also proposed a yet stronger form of the conjecture:
Conjecture 1.2.
The attractors of the replicator in any game are are exactly the content (see Definition 2.1) of the sink equilibria.
These conjectures are known to be true in some special cases of sink equilibria (Biggar and Shames, 2023b), including PNEs and attracting subgames (those where the sink equilibrium profiles are a subgame of the game). Some games, such as ordinal potential games (Monderer and Shapley, 1996) and weakly acyclic games (Young, 1993) have only PNEs as sink equilibria, so all these games satisfy the conjecture. It was subsequently proved that the stronger Conjecture 1.2 also holds for the sink equilibria of zero-sum games (Biggar and Shames, 2024).
Our contributions
We investigate the attractor-sink equilibrium hypothesis and make significant progress in understanding the possible limit behaviors of the replicator dynamic. First, in Section 2 we show through a topological argument and construction that both Conjecture 1.1 and Conjecture 1.2 fail in general games. A key component of our analysis involves understanding local sources, which are fixed points inside the sink equilibrium that are Nash equilibria of the reversed game $-u$ (Definition 2.2). Trajectories beginning near local sources move into the interior of the strategy space, which can result in paths which are not captured by the preference graph. Local sources suffice to disprove Conjecture 1.2, and for Conjecture 1.1 they suffice to disprove it in three-player games.
Disproving Conjecture 1.1 in two-player games requires a more complex argument. In Section 2.2 we analyze a particular $2\times 3$ game (Figure 4(b)) in detail, establishing the existence of a trajectory between two fixed points whose support is each $2\times 2$. We use this game as a gadget, and by composing several copies of it we are able to arrive at a two-player game with two sink equilibria, but only one replicator attractor, disproving Conjecture 1.1.
Finally, we use the insights gained from these counterexamples to identify a simple local property called pseudoconvexity, which suffices to ensure that the content of a sink equilibrium is an attractor (so Conjecture 1.2 holds). Pseudoconvexity is a property of the $2\times 2$ subgames which intersect a sink equilibrium, and so can be easily established computationally. Roughly speaking, pseudoconvexity requires that the net inflow to the sink equilibrium in the neighborhood of ‘concave’ $2\times 2$ subgames is positive. This property generalizes all the previous classes of games where Conjecture 1.2 was known to hold: potential games (Monderer and Shapley, 1996), games where the sink equilibrium is a subgame (Biggar and Shames, 2023b) and zero-sum games (Biggar and Shames, 2024). Remarkably, we show that pseudoconvexity is sufficient to ensure that no local sources can exist, and further that a pseudoconvex sink equilibrium always completely defines an attractor. This is a useful step towards a full characterization of the attractors of the replicator dynamic for two-player games.
Overall we find that while the attractor–sink equilibrium hypothesis fails in general, a broad correspondence between attractors and the preference graph does seem to hold. In Section 4 we discuss open problems inspired by our work, and hypothesize how further developments of these combinatorial tools can continue to build our understanding of learning in games.
Related work
Learning in games has a long and complex history (Cesa-Bianchi and Lugosi, 2006). In this paper we focus on the replicator dynamic (Taylor and Jonker, 1978), the flagship dynamic in this field. The replicator dynamic arose from the work of Maynard Smith on evolutionary game theory (Smith and Price, 1973), being named and formalized in (Taylor and Jonker, 1978). Since then, it has retained its central role in evolutionary game theory (Sandholm, 2010; Hofbauer and Sigmund, 2003) as well as online learning, where it is the continuous-time analogue of the multiplicative weight method (Arora et al., 2012). Finding its attractors is a central goal of the study of the replicator, both in evolutionary game theory (Zeeman, 1980; Sandholm, 2010) and more recently in learning (Papadimitriou and Piliouras, 2016, 2018). Since the work of Papadimitriou and Piliouras (2019), a line of research has developed relating the replicator and the sink equilibria. Omidshafiei et al. (2019) used the sink equilibria as an approximation of the outcome of games for the purpose of ranking the strength of game-playing algorithms. Similarly, Omidshafiei et al. (2020) used the preference graph as a tool for representing the space of games for the purposes of learning. Later, Biggar and Shames wrote a series of papers on the preference graph and its relationship to the attractors of the replicator dynamic (Biggar and Shames, 2023a, b, 2024). Another recent work (Hakim et al., 2024) explored the problem of computing the limit distributions over sink equilibria, given some prior over strategies. Our work extends the frontier of this line of investigation.
Preliminaries
We study normal-form games with a finite number $N$ of players and finitely many strategies $S_{1},S_{2},\dots,S_{N}$ for each player (Myerson, 1997). The payoffs in the game are defined by a utility function $u:\prod_{i=1}^{N}S_{i}\to^{N}$. We call a game with $|S_{1}|=m_{1},|S_{2}|=m_{2},\dots,|S_{N}|=m_{N}$ an $m_{1}\times m_{2}\times\dots\times m_{N}$ game. For simplicity, we define $Z:=\prod_{i=1}^{N}S_{i}$ as the set strategy profiles (an $N$-tuple consisting of a strategy for each player). We use $p_{-i}$ to denote an assignment of strategies to all players other than $i$. This allows us to write a strategy profile $p$ as a combination of a strategy $s\in S_{i}$ for player $i$ and the remaining strategies $p_{-i}$, which we denote by $(s;p_{-i})$. A subgame is the game resulting from restricting each player to a subset of their strategies.
A mixed strategy is a distribution over a player’s pure strategies, and a mixed profile is an assignment of a mixed strategy to each player. We also refer to strategies as pure strategies and profiles as pure profiles to distinguish them the mixed cases. For a mixed profile $x$, we write $x^{i}$ for the distribution over player $i$’s strategies, and $x^{i}_{s}$ for the $s$-entry of player $i$’s distribution, where $s\in S_{i}$. The strategy space of the game is the set of mixed profiles, which is given by $\prod_{i=1}^{N}\Delta_{|S_{i}|}$ where $\Delta_{|S_{i}|}$ are the simplices in ${}^{|S_{i}|}$. We denote $\prod_{i=1}^{N}\Delta_{|S_{i}|}$ by $X$. This is the mixed analogue of $Z$.
Any mixed profile $x$ defines a product distribution over profiles, which we denote by
$z=x^{1}\otimes x^{2}\otimes\dots\otimes z^{N}.$ (1)
That is, if $p=(s_{1},s_{2},\dots,s_{N})$ is a strategy profile, and $x$ is a mixed strategy, then $z_{p}=x^{1}_{s_{1}}x^{2}_{s_{2}}\dots x^{N}_{s_{N}}$ is the $p$-entry of the distribution $z$.
The utility function can be naturally extended to mixed profiles, by taking the expectation over strategies. We denote this by $\mathbb{U}:X\to^{N}$, defined by
$\mathbb{U}_{i}(x)=\sum_{s_{1}\in S_{1}}\sum_{s_{2}\in S_{2}}\dots\sum_{s_{N}\in S_{N}}\left(\prod_{j=1}^{N}x^{j}_{s_{j}}\right)u_{i}(s_{1},s_{2},\dots,s_{N})$ (2)
A Nash equilibrium of a game is a mixed profile $x$ where no player can increase their expected payoff by a unilateral deviation of strategy. More precisely, a Nash equilibrium is a point where for each player $i$ and strategy $s\in S_{i}$,
$\mathbb{U}_{i}(x)\geq\mathbb{U}_{i}(s;x_{-i})$ (3)
A Nash equilibrium is pure (a PNE) if $x$ is a pure profile.
Much of this paper concerns two-player games. In this context we often use much simper notation. We define an $n\times m$ game by a pair of $n\times m$ matrices $(A,B)$ where $A_{i,j}$ defines the payoff to the first player in the profile $(i,j)$. We use $x$ and $y$ to denote mixed strategies for the first and second player respectively, so their expected payoffs are $Ay$ and $x^{T}B$ respectively.
The replicator dynamic is a continuous-time dynamical system (Sandholm, 2010; Hofbauer and Sigmund, 2003) defined as the solution of the following ordinary differential equation.
$\dot{x}_{s}^{i}=x_{s}^{i}\left(\mathbb{U}_{i}(s;x_{-i})-\sum_{r\in S_{i}}x_{r}^{i}\mathbb{U}_{i}(r;x_{-i})\right),$
where $x$ is a mixed profile, $i$ a player and $s$ a strategy. It is known to be the limit of the multiplicative weights update algorithm (Arora et al., 2012) when the time step goes to zero.
A central notion in dynamical systems is the concept of an attractor (Strogatz, 2018). First, fix a dynamic. An attracting set under that dynamic is a set $S$ of points with these two properties: (1) there is a neighborhood $N\supset S,N\neq S$ that is closed under the dynamic (if the dynamic starts in $N$ it will stay there), and all points of $N$ converge to $S$ under the dynamic. An attracting set is an attractor if it is minimal, that is, no proper subset of it is attracting. We will similarly need the notion of the forward and backward limit sets. Given a point $x$, its (forward) limit set $\omega(x)$ under a given dynamic is the set of accumulation points of the orbit starting at $x$, which can be defined as
$\omega(x)=\cap_{t\geq 0}\overline{\{x(s):s\geq t\}}$
Similarly, the backward limit set $\alpha(x)$ of $x$ is the set of accumulation points of the orbit proceeding backwards in time.
Preference graphs
The preference graph of a game (Biggar and Shames, 2023a) is a directed graph whose nodes are the profiles of the game. Two profiles are $i$-comparable if they differ in the strategy of player $i$ only, and they are comparable if they are $i$-comparable for some $i$. There is an arc between two profiles if they are comparable, and the arc is directed towards the profile where that player receives higher payoff.
[…]
$W_{q,p}:=u_{i}(q)-u_{i}(p)$ (4)
[…]
Finally, the sink equilibria of a game are the sink strongly connected components of the preference graph of the game. For a recent summary of preference graph results, see Biggar and Shames (2025).
2 Refuting the Conjectures
As we discussed in the introduction, the attractor–sink equilibrium hypothesis originates with Papadimitriou and Piliouras (2019) who proposed the sink equilibria as a proxy for attractors. From a technical perspective, the first connection between these concepts was shown by Biggar and Shames (2023b) who proved that every attractor of the replicator dynamic contains the content (Def. 2.1) of some sink equilibrium. This is our starting point for this paper.
Definition 2.1 (Biggar and Shames (2023b)).
Let $H$ be a set of profiles in a game. The content of $H$, denoted $\operatorname{content}(H)$, is the set of all mixed profiles $x$ where all profiles in the support of $z:=\bigotimes_{i}x^{i}$ (Eq. (1)) are in $H$.
In particular, this implies that the attractors of the replicator dynamic are finite in number (this was not known before). Moreover, this result is not too specific to the replicator. It relies only on two key properties of the replicator: volume conservation and subgame-independence (Hofbauer and Sigmund, 1998; Sandholm, 2010). Volume conservation prohibits any asymptotically stable set from being in the interior of the strategy space. Subgame-independence asserts that (1) each subgame is invariant under the dynamic and (2) the dynamic is unaffected by strategies outside the current subgame. More generally, both of these properties extend to dynamics beyond the replicator, such as the steep FTRL dynamics of (Vlatakis-Gkaragkounis et al., 2020).
Theorem 5.2 of Biggar and Shames (2023b) proved that sink equilibria are contained in the attractors of the replicator. However, Conjecture 1.2 goes further, stating that the content of the sink equilibrium characterizes the attractors exactly. In the next subsection we provide the first refutation of this conjecture in general games by constructing a game with where the content of the sink equilibrium is not an attractor. We also refute Conjecture 1.1, the more general the ‘one-to-one’ form of the conjecture.
Our counterexamples make use of a common building block, called a local source of a sink equilibrium. As a concrete example, we will consider the game in Figure 2.
[…]
figs/coggame
[…]
figs/corner_variant
Figure 2: A preference graph (Fig. 2(a)) whose sink equilibrium (highlighted in gray) has a local source $a$ (Def. 2.2). The point $\hat{x}$ represents the interior Nash equilibrium of the $2\times 2$ subgame in the top left, shown separately in Fig. 2(b). The presence of the local source at $a$ implies that any replicator attractor of the game which contains $a$ must also contain $\hat{x}$.
This game has a unique sink equilibrium $H$ (highlighted in gray), which demonstrates an interesting property: despite being globally the unique sink connected component of the graph, $H$ is not a sink in all subgames. In particular, the node $a$ is inside $H$, but is (in one $2\times 2$ subgame, Fig. 2(b)) a source. We call the profile $a$ a local source of $H$ (Def. 2.2).
Definition 2.2 (Local sources of sink equilibria).
Let $u$ be an $N$-player game, with $x$ be a mixed profile, $H$ a sink equilibrium, and $Y$ a subgame. Then $x$ is a local source of $H$ in $Y$ if, (1) $x\in\operatorname{content}(H)\cap Y$, (2) $Y\not\subseteq\operatorname{content}(H)$ and (3) $x$ is a quasi-strict Nash equilibrium of the negated game $-u$ restricted to $Y$. We will call a point $x$ a local source of $H$ if there exists a subgame $Y$ such that $x$ is a local source of $H$ in $Y$.
In English, this definition can be read as: a local source $x$ of $H$ in $Y$ is a mixed profile in $Y$ which is contained in the sink equilibrium $\operatorname{content}(H)$, yet ‘locally’ (in the subgame $Y$) looks like a source, in the sense that all players can strictly improve their payoff by deviating to any currently unplayed strategy in $Y$. This is the opposite requirement to that of $x$ being a quasi-strict Nash equilibrium of $Y$, and so this is equivalent to $x$ being a quasi-strict Nash equilibrium of $-u$ in $Y$ [^1] Quasi-strict equilibria are those where, the payoff for deviation to any strategy outside the support of the equilibrium is strictly worse. Nash equilibria are generically quasi-strict (Harsanyi, 1973).. As an example, we return to Figure 2. Let $Y$ be the $2\times 2$ subgame shown in Fig. 2(b). The sink equilibrium $H$ intersects $Y$, but $Y\not\subseteq\operatorname{content}(H)$. The profile $a$ is in $\operatorname{content}(H)\cap Y$, and in this subgame $a$ is a source (Fig. 2(b)), so $a$ is a PNE of this $2\times 2$ subgame in the negated game $-u$. Hence $a$ is a local source of $H$.
Note that the second requirement ($Y\not\subseteq\operatorname{content}(H)$) is important. Without it, any pure profile in a sink equilibrium that is not a PNE would be a local source, because it is a source in the $1\times 2$ subgame defined by an outgoing arc of the preference graph. The notion of a local source only should include those where the associated subgame is not already accounted for inside $\operatorname{content}(H)$.
Many sink equilibria have no local sources. For instance, it is straightforward to check that a sink equilibrium that is a subgame has no local sources. Similarly, sink equilibria of zero-sum games also do not have local sources. We will show in the next section that these sink equilibria actually satisfy a stronger requirement we call pseudoconvexity, which implies that no local sources exist.
Armed with the definition of a local source, we return to Conjecture 1.2. In Figure 2, the node $a$ is a local source of $H$ in the subgame $Y$ (Fig. 2(b)). This subgame is a $2\times 2$ coordination game, where we know a trajectory exists from each source (in this case $a$) to the Nash equilibrium in the interior of this subgame ($\hat{x}$) (Hofbauer and Sigmund, 1998) [^2] When we say a trajectory exists between two fixed points $x$ and $y$, we do not mean that a trajectory beginning at $x$ will reach $y$. This is because $x$ is fixed — trajectories beginning at $x$ stay there forever. We mean rather that there are points arbitrarily close to $x$ which converge to $y$. More formally, there is a trajectory from $x$ to $y$ if there is a point $z$ where $\alpha(z)=\{x\}$ and $\omega(z)=\{y\}$.. However $\hat{x}$ is not in the content of the sink equilibrium, because its support includes the profile $p$. The existence of this trajectory implies that any attracting set containing $a$ must also contain $\hat{x}$, contradicting Conjecture 1.2.
This example is one instance of a more general fact: any sink equilibrium possessing a local source causes Conjecture 1.2 to fail.
Lemma 2.3.
If $H$ is a sink equilibrium, and $H$ has a local source, then the $\operatorname{content}(H)$ is not an attractor.
Proof.
Let $\hat{x}$ be a local source of $H$ in a subgame $Y$. Because $Y\not\subseteq\operatorname{content}(H)$, we know that $\operatorname{content}(H)\cap\operatorname{int}Y=\emptyset$. Thus we only need to show that at there is at least one point $z\in\operatorname{int}Y$ with $\alpha(z)=\{\hat{x}\}$. Because $\hat{x}$ is a quasi-strict Nash equilibrium of $-u$, this is equivalent to showing that, in $-u$, there is a $z\in\operatorname{int}Y$ with $\omega(z)=\{\hat{x}\}$. We prove this claim in Lemma A.1, by observing that the stable manifold of $\hat{x}$ in $-u$ must necessarily intersect the interior of $X$. ∎
Local sources also lead to the refutation of Conjecture 1.1, but more complex arguments are required. We will first demonstrate it in game with at least three players; the two-player case is more subtle, and we defer it to Section 2.2.
2.1 Games with three players
[…]
figs/3player_counter
[…]
figs/3player_counter_small
Figure 3: A 3-player counterexample. We show that a replicator trajectory exists from $a$ to $\hat{x}$ (a fully-mixed Nash equilibrium of the subgame in 3(b)) and also from $\hat{x}$ to $b$, implying that any attractor containing $a$ must also contain $b$.
Consider a game which possessing the preference graph in Figure 3(a). This game has three players with 3, 3 and 2 strategies respectively. Its preference graph has two sink equilibria $H_{a}$ and $H_{b}$, which we have each highlighted in gray. We have named two nodes $a$ and $b$, the former in $H_{a}$ and the latter in $H_{b}$. The critical features of the example lie in one $2\times 2\times 2$ subgame, which we isolate in Figure 3(b). The remainder of the graph serves to ensure that $H_{a}$ is a sink equilibrium of the game.
In this subgame, node $a$ is a source and node $b$ is a sink. That is, $a$ is again a local source of $H_{a}$! Because $b$ is in a different sink equilibrium to $a$, there are necessarily no paths from $a$ to $b$ in this subgraph. However, it is not possible for an attractor to contain $a$ and not $b$.
Lemma 2.4.
In Figure 3, any attracting set of the replicator containing $a$ must contain $b$.
Proof.
We will focus on the $2\times 2\times 2$ subgame in Figure 3(b) (recall that the replicator is subgame-independent). For simplicity we will assume that the weight on every arc in this subgame has value one. Each player has exactly two strategies in this subgame, and we will change make a change of variables to represent these by a single variable each. From here, $x_{1},x_{2},x_{3}$ represent the proportion that each player plays one of their strategies, such that $a=(0,0,0)$ and $b=(1,1,1)$. This subgame also contains a Nash equilibrium at $\hat{x}=(0.5,0.5,0.5)$, which is the only fixed point of the replicator in the interior of this subgame.
Because of the unit payoffs, a simple expansion of the replicator equation on this subgame gives us:
$\dot{x}_{1}=x_{1}(1-x_{1})\left((1-x_{2})(1-x_{3})-x_{2}(1-x_{3})-(1-x_{2})x_{3}+x_{2}x_{3}\right)$
The equations for $\dot{x}_{2}$ and $\dot{x}_{3}$ follow by symmetry. Consider the one-dimensional diagonal subspace which contains $a$, $b$ and $\hat{x}$. This is defined by $x_{1}=x_{2}=x_{3}$. Because of symmetry, this subspace is closed—if we start in this subspace, we remain there. Hence, on this subspace we can express the behavior of the replicator by a one-dimensional dynamical system. Defining $w=x_{1}=x_{2}=x_{3}$ gives
$\dot{w}=w(1-w)\left((1-w)^{2}-w(1-w)-(1-w)w+w^{2}\right)$
This factorizes to:
$\dot{w}=w(1-w)(2w-1)^{2}$
This is always non-negative, with fixed points at 0 ($a$), $0.5$ ($\hat{x}$) and 1 ($b$). Thus there is a trajectory from any neighborhood of $a$ to $\hat{x}$, and similarly there is a trajectory from any neighborhood of $\hat{x}$ to $b$. Any attracting set containing $a$ must also contain $\hat{x}$, and hence also $b$. ∎
2.2 Games with two players
The previous technique does not work in two-player games. The reason is that every two-player game containing a source and a sink necessarily has a path in the preference graph from the source to the sink. In the example above, our construction used a subgame (Figure 3(b)) containing both a source ($a$) and sink ($b$) but with no path between them.
It turns out that the conjecture still fails in two-player games, though the argument is more subtle. We will use the graph in Figure 4.
[…]
figs/2player_counter
[…]
figs/2player_small
Figure 4: A two-player counterexample. We show that replicator trajectories exist from $a$ to $\hat{x}$ to $\hat{y}$ to $c$ to $b$, meaning that $b$ must be in any attractor containing $a$, despite it being in a different sink equilibrium.
Like before, we have a game containing two nodes $a$ and $b$, where there are no paths from $a$ to $b$ in the graph. The node $a$ is contained in a sink equilibrium $H_{a}$ and $b$ is contained in a different sink equilibrium $H_{b}$. We have highlighted these in gray. For clarity we have omitted some arcs from the graph, where they are implied by the fact that $H_{a}$ and $H_{b}$ are sink equilibria.
Despite the apparent complexity, the key step can be reduced to reasoning about local sources in a much smaller subgame, shown in Figure 4(b). The points $\hat{x}$, $\hat{y}$ and $\hat{z}$ are fixed points of the replicator dynamic in $2\times 2$ subgames — that is, they are Nash equilibria in their respective subgames. First, consider the $2\times 2$ subgame containing $a$ and $\hat{x}$. The profile $a$ is a local source of this subgame, and just as we argued in the previous subsection, there is a trajectory from $a$ to $\hat{x}$, implying that any attracting set contain $a$ must contain $\hat{x}$. We will now show that, for some choices of arc weights, $\hat{x}$ is itself a local source of the $2\times 3$ subgame depicted in Fig. 4(b).
Lemma 2.5.
Let $u$ be a generic $2\times 3$ game whose preference graph is isomorphic to that in Figure 4(b). There exist two fixed points $\hat{x}$ and $\hat{y}$ whose support is $2\times 2$, and exactly one of these is a Nash equilibrium. There is a trajectory between these points, beginning at the non-Nash fixed point and ending at the Nash equilibrium.
Proof.
We prove this lemma in the appendix, by showing that the set of points $z$ with $\omega(z)=\{\hat{y}\}$ and $\alpha(z)=\{\hat{x}\}$ must intersect in the interior of the strategy space. ∎
Using this lemma, we return to Figure 4. By choosing appropriate weights, we can make $\hat{y}$ be a Nash equilibrium in the $2\times 3$ subgame containing $\hat{x}$ and $\hat{y}$, and make $\hat{z}$ a Nash equilibrium in the $2\times 3$ subgame containing $\hat{y}$ and $\hat{z}$. By Lemma 2.5, there is a trajectory from $\hat{x}$ to $\hat{y}$ and similarly there is a trajectory from $\hat{y}$ to $\hat{z}$. By the same argument as the case of $a$ and $\hat{x}$, there is a trajectory from $\hat{z}$ to $c$. Finally, we complete the argument by observing that these trajectories form a heteroclinic orbit from $a$ to $\hat{x}$ to $\hat{y}$ to $\hat{z}$ to $c$ and finally to $b$. This implies that any attracting set which contains $a$ must necessarily contain $b$. Each attractor contains at least one sink equilibrium, and distinct attractors are disjoint. Because any attractor containing $H_{a}$ must also contain $H_{b}$, we conclude that this game has only a single attractor, despite having two sink equilibria.
3 Pseudoconvex sink equilibria are attractors
Our understanding of stability under the replicator dynamic is becoming clearer. When the sink equilibrium has a very simple structure, such as a subgame, its content is an attractor. PNEs are a special case of this. These sinks have no local sources. Sink equilibria of zero-sum games can have non-trivial structure, albeit subject to significant constraints, and they too define attractors (Biggar and Shames, 2024). On the other hand, the counterexamples above demonstrate that the presence of local sources can cause some sink equilibria to not be attractors. What special properties of the sink equilibria in these games prevent the occurrence of local sources, and ensure stability? To better understand this, we will examine the structure of their $2\times 2$ subgames. In a $2\times 2$ subgame, the presence of a local source requires that exactly three of the profiles are within the sink equilibrium. We call such cases a cavity of the sink equilibrium.
Definition 3.1.
Let $Y$ be a $2\times 2$ subgame of a game. If exactly three of the profiles in $Y$ are contained in a sink equilibrium $H$, we call this subgame a cavity of $H$.
[…]
figs/corner
[…]
figs/corner2
[…]
figs/corner3
Figure 5: A cavity (Def. 3.1) of a sink equilibrium $H$, where $x\not\in H$ and $w,y,z\in H$. Because $x\not\in H$, the arcs at $x$ are necessarily directed towards $y$ and $z$ respectively. Up to symmetry, there are three cases, shown in Figs. 5(a)5(b) and 5(c).
Observe that a sink equilibrium $H$ is a subgame if and only if it has no cavities. Equivalently, $H$ is a subgame if and only if $\operatorname{content}(H)$ is convex. Cavities are $2\times 2$ subgames where the sink equilibrium is locally concave. The absence of cavities is sufficient to guarantee that $\operatorname{content}(H)$ is an attractor, but it is a very strong requirement, because it implies the sink equilibrium is a subgame. Can a weaker requirement suffice? Examining Figure 5, we note that case (c) is a local source and so clearly will be a problem, but case (b) is a local sink — all points converge uniformly to the sink equilibrium. Finally, in case (a), all points approach the sink equilibrium but may or may not do so uniformly, depending on the relative sizes of $b$ and $a$.
Definition 3.2.
Let $w=(\alpha,\beta)$ be a profile in a sink equilibrium $H$, and $x=(\gamma,\delta)$ be another profile outside $H$ such that the subgame $Y=\{\alpha,\gamma\}\times\{\beta,\delta\}$ is a cavity of $H$. We say $Y$ is pseudoconvex if $W_{(\gamma,\beta),(\alpha,\beta)}+W_{(\alpha,\delta),(\alpha,\beta)}<0$.
That is, $Y$ is locally pseudoconvex if the concavity of $Y$ is ‘not too severe’ — the sum of weights of the arcs into the profile $w$ is positive. Using the language of Figure 5, $Y$ is pseudoconvex if either it is type 5(a), or it is type 5(a) and $a\leq b$. This prevents local sources (type 5(c)) in $2\times 2$ subgames. However, quite remarkably, this is also sufficient to guarantee that the entirety of $\operatorname{content}(H)$ is an attractor (Theorem 3.7).
Definition 3.3 (Pseudoconvex sink equilibria).
A sink equilibrium is pseudoconvex if every cavity is pseudoconvex.
It is clear that if a sink equilibrium is a subgame it is trivially pseudoconvex. It is also true — though much less obvious — that the sink equilibria of zero-sum games are pseudoconvex.
Lemma 3.4.
The sink equilibria of two-player zero-sum games are pseudoconvex.
Proof.
[…]
Hence for all the classes of sink equilibria where we know the content is an attractor, these sink equilibria are pseudoconvex. It is natural to therefore conjecture that pseudoconvexity is sufficient for the content to not be an attractor. This turns out to be true for two-player games, as we now show (Theorem 3.7). To prove this we will need some novel formalisms and lemmas.
Definition 3.5 (Product matrix).
Let $u$ be a $m_{1}\times m_{2}\times\dots\times m_{N}$ game. The product matrix of $u$ is the following matrix $M\in^{(\prod_{i}m_{i})\times(\prod_{i}m_{i})}$, indexed by profiles $p$ and $q$ in $Z$:
$M_{q,p}=\sum_{i=1}^{N}\left(u_{i}(q_{i};p_{-i})-u_{i}(p)\right)=\sum_{i=1}^{N}W_{(q_{i};p_{-i}),p}$
Note that when $p$ and $q$ are $i$-comparable, $M_{q,p}=u_{i}(q)-u_{i}(p)=W_{q,p}$. The product matrix allows for a significantly simpler presentation of the replicator equation in the product space of the game (Lemma 3.6). This idea comes from Biggar and Shames (2024), who introduced this construction in two-player zero-sum games. Here we generalize it to $n$-player games. While straightforward, this transformation is critical for our results and we believe it is a useful fact for analyzing the replicator in other contexts as well.
Lemma 3.6.
Let $u$ be a $N$-player game, $M$ the product matrix of $u$ and $p=(s_{1},s_{2},\dots,s_{N})$ a profile. With $z_{p}:=\prod_{i=1}^{N}x^{i}_{s_{i}}$ (Eq. (1)), then under the replicator dynamic:
$\dot{z}_{p}=z_{p}(Mz)_{p}$
Proof.
By Lemma A.1 of Biggar and Shames (2023b), the replicator equation can be written
$\dot{x}^{i}_{s}$ $=x_{s}^{i}\sum_{r\in S_{i}}x_{r}^{i}\sum_{p_{-i}\in Z_{-i}}z_{p_{-i}}\left(u_{i}(s;p_{-i})-u_{i}(r;p_{-i})\right)$
$=x_{s}^{i}\sum_{p_{-i}\in Z_{-i}}\sum_{r\in S_{i}}x_{r}^{i}z_{p_{-i}}\left(u_{i}(s;p_{-i})-u_{i}(r;p_{-i})\right)$
$=x_{s}^{i}\sum_{p_{-i}\in Z_{-i}}\sum_{r\in S_{i}}z_{(r;p_{-i})}\left(u_{i}(s;p_{-i})-u_{i}(r;p_{-i})\right)$
$=x_{s}^{i}\sum_{p\in Z}z_{p}\left(u_{i}(s;p_{-i})-u_{i}(p)\right)$
Now we observe that for $q=(q_{1},q_{2},\dots,q_{N})$,
[…]
$=z_{q}\sum_{p\in Z}z_{p}\sum_{i=1}^{N}\left(u_{i}(q_{i};\bar{p})-u_{i}(p)\right)=z_{q}\sum_{p\in Z}z_{p}M_{q,p}$
$=z_{q}(Mz)_{q}$
Theorem 3.7.
If a sink equilibrium of a two-player game is pseudoconvex, then its content is an attractor of the replicator.
Proof.
Let $H$ be a sink equilibrium, which we assume is pseudoconvex. We will show it is asymptotically stable by a Lyapunov argument, using a similar structure to Theorem 4.3 of Biggar and Shames (2024). First, we define $z_{H}:=\sum_{h\in H}z_{h}$. That is, $z_{H}$ is the cumulative total distributed over the profiles in $H$ in the product distribution $z$ (Eq. 1). Note that $z_{H}=1$ if and only if $x\in\operatorname{content}(H)$. $z_{H}$ is uniformly continuous, and so to prove that $\operatorname{content}(H)$ is an attractor it is sufficient to show that $\dot{z}_{H}>0$ in some neighborhood of $\operatorname{content}(H)$. Now fix some small $01>\epsilon>$ 0. We will assume that $z_{H}=1-\epsilon$, and we will show that for small enough $\epsilon$, $\dot{z}_{H}>0$.
From Lemma 3.6 we have that $\dot{z}_{H}=\sum_{h\in H}z_{h}(Mz)_{h}=\sum_{p\in Z}\sum_{h\in H}z_{p}z_{h}M_{h,p}$. Each term in this sum corresponds to a pair of profiles $p$ and $h$ with $h\in H$. First, we divide this sum into comparable and non-comparable pairs of profiles:
$\dot{z}_{H}=\sum_{p,h\in H\text{ comparable}}z_{p}z_{h}M_{h,p}+\sum_{p,h\in H\text{ not comparable}}z_{p}z_{h}M_{h,p}$
If $p$ and $h$ are comparable and $p\in H$, then the first sum contains the terms $z_{p}z_{h}M_{p,h}$ and $z_{p}z_{h}M_{h,p}$. Because they are comparable, $M_{p,h}=W_{p,h}=-W_{h,p}=-M_{h,p}$ (Definition 1), so these terms cancel. The sum becomes
$\dot{x}_{H}=\sum_{p\not\in H,h\in H\text{ comparable}}z_{p}z_{h}M_{h,p}+\sum_{p,h\in H\text{ not comparable}}z_{p}z_{h}M_{h,p}$
[…]
Now consider the case where $p$ and $h$ are non-comparable. Suppose $p=(\alpha,\beta)$ and $h=(\gamma,\delta)$ where none of these are equal, because $p$ and $h$ are not comparable. Let $a:=(\alpha,\delta)$ and $b:=(\gamma,\beta)$ be the other two profiles in this $2\times 2$ subgame. First observe that $z_{p}z_{h}=x_{\alpha}y_{\beta}x_{\gamma}y_{\delta}=z_{a}z_{b}$. So, for instance, if $z_{p}z_{h}M_{h,p}$ and $z_{a}z_{b}M_{a,b}$ are both in the sum, we can combine them into a single term $z_{p}z_{h}(M_{h,p}+M_{a,b})$. We will now group all such terms with common $z$-coefficients. Which of these pairs actually appear in the sum depends on which of $p,a,b,h$ are in $H$. We know $h\in H$, so the remaining cases are:
  1. 1.
    All of $p,a,b,h$ are in $H$: The sum contains all of the terms $M_{p,h}+M_{h,p}+M_{a,b}+M_{b,a}$. Expanding this by Definition 1 gives $W_{a,p}+W_{b,p}+W_{a,h}+W_{b,h}+W_{p,b}+W_{h,b}+W_{h,a}+W_{p,a}=0$, because each term $W_{i,j}=-W_{j,i}$. Hence if all are in $H$, these terms cancel in the sum.
  2. 2.
    Three are in $H$ — we assume w.l.o.g. that $p\not\in H$ and $a,b,h\in H$. This is a cavity of $H$. Then the terms in the sum are $M_{h,p}+M_{a,b}+M_{b,a}$. By the same argument as above, $M_{h,p}+M_{a,b}+M_{b,a}=-M_{p,h}$, which is equal to $M_{p,h}=W_{a,h}+W_{b,h}$. Hence, $M_{p,h}$ is positive if and only if this cavity is pseudoconvex. By assumption, all cavities are pseudoconvex, so $-W_{p,h}=M_{h,p}+M_{a,b}+M_{b,a}$ is non-negative.
  3. 3.
[…]
  4. 4.
    Only $h$ is in $H$. This last case is the most difficult, because here the terms can be negative. Each such pair contains only the term $z_{p}z_{h}M_{h,p}$. However, note that because $a,b\not\in H$, $z_{a},z_{b}<\epsilon$. Hence $z_{p}z_{h}=z_{a}z_{b}<\epsilon^{2}$. Thus all negative terms in $\dot{x}_{H}$ have coefficient at most $\epsilon^{2}$.
We have now grouped the terms in this sum by their distinct $z$-coefficients, so each term has the form $x_{\alpha}y_{\beta}x_{\gamma}y_{\delta}(M_{i,j}+\dots)$. For simplicity, we write $K_{\alpha,\beta,\gamma,\delta}:=\sum M_{i,j}$ where $z_{i}z_{j}=x_{\alpha}y_{\beta}x_{\gamma}y_{\delta}$. We now define $\mu:=\min\{K_{\alpha,\beta,\gamma,\delta}\ :\ K_{\alpha,\beta,\gamma,\delta}>0\}$ and similarly $m:=\max\{|K_{\alpha,\beta,\gamma,\delta}|\ :\ K_{\alpha,\beta,\gamma,\delta}<0\}$. These are the smallest and largest positive and negative terms respectively. There are at most $N^{2}$ terms in this sum, where $N$ is the number of profiles. By the above, each negative term has coefficient at most $\epsilon^{2}$, so the total sum of negative terms is at most $-mN^{2}\epsilon^{2}$. Now select an $h\in H$ where $z_{h}\geq(1-\epsilon)/|H|$. Since $z_{H}=1-\epsilon$, such a node must exist. Then:
$\dot{x}_{H}\geq\sum_{p\not\in H,\ z_{p}z_{h}=x_{\alpha}y_{\beta}x_{\gamma}y_{\delta},\ K_{\alpha,\beta,\gamma,\delta}>0}z_{p}z_{h}\mu-mN^{2}\epsilon^{2}=\mu z_{h}\sum_{p\not\in H,\ z_{p}z_{h}=x_{\alpha}y_{\beta}x_{\gamma}y_{\delta},\ K_{\alpha,\beta,\gamma,\delta}>0}z_{p}-mN^{2}\epsilon^{2}$
This inequality holds because we have retained the contribution from all negative terms (in the $mN^{2}\epsilon^{2}$ term), and reduced the set of positive terms. Specifically we have included terms where the coefficient equals $z_{p}z_{h}$ for our fixed $h$ and where $p\not\in H$. Now we must determine the sum $\sum z_{p}$ over these $p$. The total sum over $z_{p}$ with $p\not\in H$ is $\epsilon$, but some $z_{p}$ are not included in this sum, if they correspond to a negative term $z_{w}z_{h}K_{\alpha,\beta,\gamma,\delta}<0$. However, by the argument above, this occurs only in case (4). There, the remaining profiles $a$ and $b$ are not in $H$, and so contribute two positive terms $z_{a}z_{h}W_{h,a}$ and $z_{b}z_{h}W_{h,b}$ to this sum. By this argument, at least 2/3 of the terms in this sum must be positive. Also, each such $z_{p}$ has $z_{p}<\epsilon^{2}$. We obtain
$\dot{z}_{H}$ $\geq k((1-\epsilon)/|H|)(\epsilon-(1-|H|)\epsilon^{2}/3)-mN^{2}\epsilon^{2}$
$\geq k\epsilon/|H|-o(\epsilon^{2})$
Thus, for small enough $\epsilon>0$, this term is strictly positive. ∎
We believe that this result is a major step towards our ultimate goal: a polynomial-time algorithm which, given a two-player game in normal form, identifies its attractors. When the sink equilibria are pseudoconvex, we know the attractors: they are the content of these sink equilibria. But if a local source exists in a sink component, some replicator paths “escape” the component. However finding local sources, or determining if any exist, can be very hard. The power of this theorem is that pseudoconvexity (a $2\times 2$ property) is sufficient to guarantee, non-constructively, that no such escaping trajectories exist. If the sink equilibrium is not pseudoconvex, the proof of the theorem does not seem to provide guidance on where to look for these escaping trajectories. Analyzing the attractors of two-player games beyond pseudoconvexity is an important open research problem left by this work.
4 Conclusions and open problems
Let us return to our original goal: to understand, and ultimately compute, the possible long-run outcomes of game dynamics by exploiting their relationship to sink equilibria. We have made some significant progress on this problem for the replicator dynamic. For games where the sink equilibria are pseudoconvex, we have obtained a precise and efficiently-computable characterization of the attractors. On the other hand, we have also shown how this simple picture does not apply to all games. The presence of local sources can lead to attractors that are larger than the sink equilibria, sometimes merging two or more sink equilibria. However, local sources appear to have a certain combinatorial structure, and we believe that further investigation will reveal more insight into the structure of replicator attractors.
This work opens up a number of important open problems.
Local sources and paths.
We know that the presence of local sources can lead to trajectories which escape sink equilibria. Yet local sources too possess a specific graph-theoretic structure — we believe that a broader combinatorial framework generalizing the preference graph could incorporate this case, and thus potentially characterize the ultimate structure of the attractors of a game.
Beyond uniform convergence.
The criteria for a set to be an attractor are very strong — in particular they require a neighborhood of the set to exist where all trajectories approach the attractor. However, for the replicator dynamic, there exist cases where this requirement seems too demanding. Figure 6 shows a $2\times 3$ game, where the second column strategy dominates the third. Under the replicator, the third column eventually vanishes from all interior starting points, and so the dynamic converges to the highlighted $2\times 2$ subgame. But the preference graph is strongly connected, so the unique attractor is the whole game. The discrepancy is explained by the fact that the replicator does not converge uniformly to this smaller set (Biggar and Shames, 2024, 2025). This lack of uniformity follows as a consequence of subgame-independence. This suggests that stronger prediction of game dynamics could be made if we allowed for weaker notions of convergence.
[…]
figs/2x3_dominance
Figure 6: Under the replicator, all interior starting points converge to the highlighted $2\times 2$ subgame. However, this is a strict subset of the unique attractor, which is the whole strategy space (Biggar and Shames, 2025).
Algorithmic problems.
The ultimate goal of this research program is a polynomial-time algorithm which, given a game in normal form, will output the combinatorial structure of its attractors, and do this for a wide range of learning dynamics. For the case of the replicator dynamic we have made reasonable progress, with pseudoconvexity providing a simple sufficient condition for local stability. However, the possibility of local sources leads to two new problems:
  1. 1.
    We know that the presence of a local source in a sink equilibrium $H$ is sufficient to show that $\operatorname{content}(H)$ is not an attractor. But what of the converse: if a sink equilibrium is not stable, must there exist a local source acting as a ‘witness’ to this instability? From a computational perspective, given $H$, can we efficiently verify if $\operatorname{content}(H)$ is an attractor?
  2. 2.
    If a local source does exist, some additional points may need to be added to the sink equilibrium. But which? Is there an iterative procedure for adding the missing points to a sink equilibrium until we find an attractor?
Large games.
The limit behavior of learning in large multi-player games is of great interest in economics — however, we know that even the most modest algorithmic goals related to sink equilibria are PSPACE-complete for many succinct representations of games (Fabrikant and Papadimitriou, 2008). It would be very interesting to make progress in characterizing and computing the attractor structure of a game for the case of symmetric multi-player games.
Appendix A Proofs
Lemma A.1.
Let $\hat{x}$ be a quasi-strict Nash equilibrium on the boundary $\operatorname{bd}X$ of the strategy space $X$. Then there exists a point $z\in\operatorname{int}X$ with $\omega(z)=\{\hat{x}\}$.
Proof.
Let $Y$ be the subgame that is the support of $\hat{x}$, with does not include all strategies because $\hat{x}$ lies in the boundary of $X$. We begin by constructing the Jacobian of the replicator at $\hat{x}$. Let $s_{i}$ be a strategy that is not in the support of $\hat{x}^{i}$, so $\hat{x}^{i}_{s_{i}}=0$. Observe that
$\partial\dot{x}^{i}_{s_{i}}/\partial x^{i}_{s_{i}}=u_{i}(s_{i};x_{-i})-u_{i}(x)+x^{i}_{s_{i}}\partial/\partial x^{i}_{s_{i}}(u_{i}(s_{i};x_{-i})-u_{i}(x))$
and so at $\hat{x}$, where $\hat{x}^{i}_{s_{i}}=0$, we have $\partial\dot{x}^{i}_{s_{i}}/\partial x^{i}_{s_{i}}=u_{i}(s_{i};\hat{x}_{-i})-u_{i}(\hat{x})$. Similarly, for any other $x^{k}_{\ell}$,
$\partial\dot{x}^{i}_{s_{i}}/\partial x^{k}_{\ell}=x^{i}_{s_{i}}(\dots)$
so $\partial\dot{x}^{i}_{s_{i}}/\partial x^{k}_{\ell}=0$ at $\hat{x}$. Hence the strategy $s_{i}$ column of the Jacobian of the replicator at $\hat{x}$ is equal to $(u_{i}(s_{i};\hat{x}_{-i})-u_{i}(\hat{x}))\mathbf{e}_{s_{i}}$, where $\mathbf{e}_{s_{i}}$ is the standard basis vector of the $s_{i}$ column. We conclude that $\mathbf{e}_{s_{i}}$ is an eigenvector of the Jacobian at $\hat{x}$, whose eigenvalue is $u_{i}(s_{i};\hat{x}_{-i})-u_{i}(\hat{x})$ — the relative payoff of $s_{i}$ at $\hat{x}$. Because this is a quasi-strict Nash equilibrium, this is strictly negative. These are called the transversal eigenvalues of $\hat{x}$ (Hofbauer and Sigmund, 1998).
We now apply the stable manifold theorem. This tells us that, in the neighborhood of $\hat{x}$, there is an invariant manifold of points which converge to $\hat{x}$, and this manifold is tangent to the space $E^{s}$ spanned by the eigenvectors with negative eigenvalues. Because all transversal eigenvalues are negative, there is a direction in the stable manifold of $\hat{x}$ which points into the interior of the space. Hence there exists a points on the interior of $X$ which converges to $\hat{x}$. ∎
Lemma A.2 (Lemma 2.5).
Let $u$ be a generic $2\times 3$ game whose preference graph is isomorphic to that in Figure 4(b). There exist two fixed points $\hat{x}$ and $\hat{y}$ whose support is $2\times 2$, and exactly one of these is a Nash equilibrium. There is a trajectory between these points, beginning at the non-Nash fixed point and ending at the Nash equilibrium.
Proof.
(Claim: generically, exactly one of $\hat{x}$ and $\hat{y}$ is Nash.) Let $r$ be the pure strategy for player 1 which involves playing the top row. Then $\hat{x}^{1}_{r},\hat{y}^{1}_{r}\in(0,1)$ denote the probability mass on $r$ in $\hat{x}$ and $\hat{y}$ respectively. We will show that $\hat{x}$ is Nash if $\hat{x}^{1}_{r}>\hat{y}^{1}_{r}$, $\hat{y}$ is Nash if $\hat{x}^{1}_{r}<\hat{y}^{1}_{r}$ and both are Nash in the non-generic case where $\hat{x}^{1}_{r}=\hat{y}^{1}_{r}$.
Let $c_{1},c_{2},c_{3}$ be the column strategies, ordered from left to right. Given a fixed mixed strategy $(z,1-z)$ for the row player, the column player receives a payoff for each $c_{i}$ that is a linear function of $z$. At $z=0$, the column player prefers their strategies in the best-to-worst order $c_{3},c_{1},c_{2}$. At $z=1$, the column player prefers their strategies in the best-to-worst order $c_{2},c_{1},c_{3}$. At $z=\hat{x}^{1}_{r}$, the column player is indifferent between $c_{1}$ and $c_{2}$ and at $z=\hat{y}^{1}_{r}$, they are indifferent between $c_{2}$ and $c_{3}$. If $\hat{x}^{1}_{r}<\hat{y}^{1}_{r}$, then at $z=\hat{x}^{1}_{r}$ the column player must prefer $c_{3}$ over $c_{1}$ and $c_{2}$, so $\hat{x}$ is not a Nash equilibrium. By contrast, at $z=\hat{y}^{1}_{r}$, the column player prefers both $c_{2}$ and $c_{3}$ (and hence $\hat{y}^{2}$) over $c_{1}$, so $\hat{y}$ is a Nash equilibrium. The opposite is true when $\hat{x}^{1}_{r}>\hat{y}^{1}_{r}$. If $\hat{x}^{1}_{r}=\hat{y}^{1}_{r}$, then at this point the column player is indifferent between $c_{1}$, $c_{2}$ and $c_{3}$, resulting in a continuum of Nash equilibria. This completes the claim.
[…]
figs/2player_proof
Figure 7: The game from Figure 4(b), with the set $W$ highlighted as in the proof of Lemma 2.5.
(Claim: there exists a point $z$ where $\alpha(z)=\hat{x}$ and $\omega(z)=\hat{y}$.) In Figure 7 we show a more detailed depiction of this game for the purposes of the proof. The game has two sources and two sinks, which we name $r_{1}$, $r_{2}$, $s_{1}$ and $s_{2}$ respectively. The remaining two profiles are named $a$ and $d$. The sinks are PNEs, both of which are attractors of the replicator. Further, because every attractor must contain a sink equilibrium (Biggar and Shames, 2023b), these are the only attractors.
Let $\omega^{-1}(s_{1})$ and $\omega^{-1}(s_{2})$ be the sets of points $z$ whose limit set $\omega(z)$ is $\{s_{1}\}$ or $\{s_{2}\}$ respectively. These are disjoint invariant sets, and they are open by the continuity of the replicator dynamic. Hence there exists a closed set $W$ of points whose limit sets are neither $\{s_{1}\}$ nor $\{s_{2}\}$. Note that $W$ is also an invariant set. Because each proper subgame of this game is two-dimensional, we can completely characterize the intersection of $W$ and the boundary $\operatorname{bd}X$ of the strategy space by considering each subgame in turn. The result is the set depicted in Figure 7. It contains all fixed points except the sinks $s_{1}$ and $s_{2}$, as well as the one-dimensional trajectories joining them, and a two-dimensional region in the subgame spanned by $a$, $\hat{x}$, $r_{1}$ and $d$, and the trajectories between them.
By Proposition 5.1 of Hofbauer et al. (2009), the limit set of the time average of an interior point $z$ is a closed subset of $X$ that is both invariant under best-response dynamics. For $z\in\operatorname{int}X\cap W$, $\omega(z)$ is on the boundary (Hofbauer and Sigmund, 1998), and so must additionally be a subset of $W\cap\operatorname{bd}X$. The only subset of this set that is invariant under best-response dynamics is the Nash equilibrium set $\{\hat{y}\}$. Hence for any $z\in W\cap\operatorname{int}X$, $\omega(z)=\{\hat{y}\}$.
Now we repeat the steps above, but with the role of forward and backward limit sets reversed. Let $S$ be the set of points $z$ in $X$ whose backward limit set $\alpha(z)$ is neither $\{r_{1}\}$ nor $\{r_{2}\}$. Again we can characterize the set $S\cap\operatorname{bd}X$, which has an analogous structure to $W$. The flow of the replicator in reverse time is equal to the flow of the replicator on the negated game $-u$, so we can apply the same arguments about $\omega$-limits to $\alpha$-limits. We conclude, again, that for every $z\in S\cap\operatorname{int}X$, $\alpha(z)=\{\hat{x}\}$.
To complete the argument, we just need to show that $S\cap W\cap\operatorname{int}X$ is not empty. For this, observe that if $z\in\operatorname{bd}X\cap\omega^{-1}(s_{1})$, then because $\omega^{-1}(s_{1})$ is open there is an $\epsilon>0$ such that all points within $\epsilon$ of $z$ are in $\omega^{-1}(s_{1})$. In particular, there are points near $z$ in $\operatorname{int}X\cap\omega^{-1}(s_{1})$.
Now consider a neighborhood of $\hat{x}$ in $S\cap\operatorname{int}X$. $\hat{x}$ lies on the boundary of $\omega^{-1}(s_{1})$ and $\omega^{-1}(s_{2})$, so every neighborhood in $S$ contains points in both of these sets. Similarly, every neighborhood in $S\cap\operatorname{int}X$ must contain points in both of these sets, by the argument in the previous paragraph. Finally, every neighborhood of $\hat{x}$ in $S\cap\operatorname{int}X$ must therefore contain points in the boundary of $\omega^{-1}(s_{1})$ and $\omega^{-1}(s_{1})$, which is entirely contained in $W$. Hence $S\cap W\cap\operatorname{int}X$ is nonempty. ∎